OpenStudy (hba):

Given that f(x)=x^2-6x-40/x^2-16 , then f(-4)=?

OpenStudy (hba):

A) infty B) indeterminate C)7/4 or D)4/7

OpenStudy (anonymous):

Given that f(x)=x^2-6x-40/x^2-16 , then f(-4)=?. Replace every x with -4

OpenStudy (hba):

16+24-40/16-16=0/0

OpenStudy (hba):

@AbhimanyuPudi

OpenStudy (hba):

Yeah good one.@zordoloom

OpenStudy (anonymous):

There is a vertical asympote at 4. and a hole at -4.

OpenStudy (anonymous):

I think the answer is C

OpenStudy (anonymous):

Plot plot plot ...

OpenStudy (anonymous):

It can't be c.

OpenStudy (anonymous):

More like indeterminate.

OpenStudy (anonymous):

Factorise the numerator x^2-6x-40 = (x-10)(x+4) x^2-16 = (x-4)(x+4) So, f(x) = (x-10)(x+4) / (x-4)(x+4) = x-10/x-4 f(-4) = -4-10 / -4-4 = -14/-8 = 7/4

OpenStudy (anonymous):

the answer is 7/4

OpenStudy (hba):

please explain ?

OpenStudy (anonymous):

i show you now\[\lim_{x \rightarrow -4} \frac{ x ^{2} -6x-40 }{ x ^{2}-16 }\]

OpenStudy (anonymous):

@Rezz5 is correct. I totally forgot about limits.

OpenStudy (anonymous):

now use L'Hopitals rule as putting -4 in both denominator and numerator we get zero, right? su we can use L'Hopitals rule, differentiate booth top and bottom we get... \[\lim_{x \rightarrow -4} \frac{ 2x-6 }{ 2x }\] now sub in -4 in the equation, and voila, we get 7/4, if you dont understand tell me

OpenStudy (hba):

thanks a lot.

OpenStudy (anonymous):

welcome : )